### 26 Dec 2023 v1

# Large Language Model Situational Awareness Based Planning

Abstract
This work pioneers evaluating emergent planning capabilities based on situational awareness in large language models. We contribute (i) novel benchmarks and metrics for standardized assessment; (ii) a unique dataset to spur progress; and (iii) demonstrations that prompting and multi-agent schemes significantly enhance planning performance in context-sensitive planning tasks. Positioning this within a situated agent and automated planning research, we highlight inherent reliability challenges–efficiently mapping world states to actions without environmental guidance remains open despite simulated domain advances. Although out-of-scope, limitations around validation methodology and data availability indicate exciting directions, including fine-tuning on expanded planning corpora and optimizations for triggering fast latent planning. By conclusively demonstrating current methods’ promise and limitations via rigorous comparison, we catalyze investigating reliable goal-directed reasoning for situated agents.
Index Terms—  Large Language Model, Planning, Situational Awareness, Situated Agent, Multi-Agent Reasoning
1 Introduction
Developing AI agents that can make flexible decisions is challenging, as they need to handle unpredictability in the real world [1]. Humans determine appropriate interventions through situational awareness, while inadequate situation awareness has been identified as a primary cause of accidents due to human error [2, 3]. As Yadav [4] notes, rigorously examining situational awareness in large language models (LLMs) is crucial to steer their safe and reliable development. Without broader awareness, any seemingly beneficial action risks unintended harm [5, 6]. For example, an autonomous agent may need nuanced judgements to avoid potential harm in situations like a toddler reaching for a pot or playing with a knife.
Fig. 1: Large language models’ planning enhancements based on situational awareness.
Fig. 2: Iteratively generating and evaluating plans in a multi-agent proactive AI system produces iterative feedback to enhance reasoning and factual accuracy. † Refers to selecting the human demo for comparison only in the first round of iteration.
This research shows LLMs can exhibit human-like planning capacities based on situational awareness. As Fig. 1 illustrates, when prompted for situational awareness based planning spanning perception, comprehension and projection [2] or when provided reasoning feedback, the LLM displays improved deductive reasoning requiring perspective-taking and contemplating potential outcomes. This study distinguishes itself from previous AI reasoning and planning research in significant ways. Much existing work evaluates task-specific planning or makes assumptions about predefined steps under strict conditions [7, 8, 9, 10, 11, 12]. Additionally, various agents reactively generate plans only after explicit instructions or demands [13, 14, 15, 16, 17, 18, 19, 20], rather than proactively. In contrast, our approach centres on assessing and enhancing models’ capacities for proactive situational planning when faced with real, open-world dilemmas. Without rigid constraints or environmental feedback, the models must leverage deductive logic to map conjectured actions while weighing consequences using only a descriptive starting scenario and prompts. By prioritizing latent planning that mirrors human cognition - namely situational awareness based rather than domain-limited planning - our method signifies a break from former suppositions and appraisals.
The remainder of this paper overviews the methodology in Section 2, details experiments in Section 3, analyzes results in Section 4 and concludes with a summary and future directions in Section 5.
2 Methodology
In this section, we outline the key challenges and methodological components that enable collaborative multi-agent reasoning to enhance LLMs’ situational planning capacities.
2.1 Task Formulation and Key Challenges
We formulate situational awareness based planning as grounded inference over a dynamic hazard scenario s, with $s\in S$ where S denotes the hazard situation space. The input comprises an unordered set of concepts $x=\{c1,c2,...,ck\}\subseteq C$ encapsulating entities, events, and temporal evolution within s, with C the overall conceptual vocabulary. The output is a step-wise plan $\pi:S\to A$ with actions $a1,a2,...\in A$, where A denotes the action space of possible interventions. Learning the planning policy $\pi:S\to A$ requires overcoming two intrinsic challenges:
  • Failing to achieve high levels of situational awareness - lacking perception, comprehension, and projection of the hazard environment when deducing appropriate state-action mappings
  • Struggling to anticipate potential downstream consequences of planned actions on human safety and property due to inadequate understanding of hazardous situation dynamics
By formulating hazard remediation as a conceptual planning task requiring strong situational awareness, we evaluate the multidimensional latent reasoning essential for reliable situated agents operating in hazardous environments.
2.2 Multi-AI Agents Enhance Reasoning and Accuracy
Recent work has shown that employing multiple LLMs in a cooperative framework, whether collaborative or adversarial, can enhance reasoning and factuality accuracy. As Du et al. discuss, the debate between agents allows them to critique each other’s logic, correcting flaws [21]. Similarly, Liang et al. find disagreement motivates broader reasoning as agents try to outdo each other [22]. In these cases, collaboration complements individual strengths [21, 22, 23]. In this work, we use two LLM agents - $\text{LLM}_{gen}$ for plan generation and $\text{LLM}_{eval}$ for critical evaluation. We rely on the closed-loop promotion between these complementary roles to improve latent planning.
2.3 State-based Planning with feedback
Current AI systems that rely solely on rigid, context-insensitive rules risk unintended outcomes when deployed in complex, real-world environments [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. To enable more reliable and ethical decision-making, architectures should aspire to model interdependent variables and causal relationships, similar to fluid human reasoning processes.
One approach is to have LLM agents iteratively generate and assess potential solutions before implementation. As an example, we frame the design of a finite state machine (FSM) [24] as a collaborative process between two models. A latent FSM plan can be defined by a tuple $M=(S,T,A)$, comprising a set of states $S$, transitions $T$, and actions $A$. The process begins by representing the plan’s reasoning process as $R$, with a generator agent reasoning ($R_{LLM_{gen}}$) to plan then an evaluator agent reasoning ($R_{LLM_{eval}}$) to evaluate. $R_{LLM_{gen}}$ proposes a candidate FSM plan $\hat{M}$, which $R_{LLM_{eval}}$ then scores and provides feedback $f$ on. $R_{LLM_{gen}}$ incorporates this feedback into the next proposal. This iterative loop continues until the score of $\hat{M}$ is higher than that of $M^{*}$(the benchmark plan), which is finally adopted as the new optimal plan $M^{*}$. This approach, shown in Algorithm 1, allows for tight refinement loops resembling human reasoning. By assessing solutions before real-world implementation, unintended outcomes can potentially be anticipated and mitigated.
Algorithm 1 Situational Awareness Based Planning
1:  $\hat{M}\ \ \leftarrow\ {R_{LLM_{gen}}(S,T,A)}$
2:  $score$, $f$  $\leftarrow$  $R_{LLM_{eval}}{(\hat{M})}$
3:  while $\hat{M}_{score}\ <\ M^{*}_{score}$ do
4:     $\hat{M}\ \leftarrow\ {R_{LLM_{gen}}(S,T,A,f)}$
5:     $score$, $f$  $\leftarrow$  $R_{LLM_{eval}}{(\hat{M})}$
6:  end while
7:  $M^{*}\ \leftarrow\ \hat{M}$
8:  $\mathbf{adopt}\ M^{*}$
Fig. 2 visually depicts the iterative process between $\text{LLM}_{gen}$ and $\text{LLM}_{eval}$. Consider a scenario where a young child attempts to touch a hot pot on an active stove, posing a safety risk. The housekeeper robot observing this scene starts planning appropriate interventions. First, the instruction-prompted $\text{LLM}_{gen}$ leverages its reasoning capacity to imagine potential outcomes and draft candidate FSM plans to mitigate harm. For instance, abruptly obstructing the child may startle the child, suggesting a gentle approach or showing toys instead. If the child refuses intervention and gets burned, emergency responses may become necessary. $\text{LLM}_{gen}$ passes its proposed plan $\hat{M}$ with the comparison plan - which is a human demo at the first round then the previous plan - to $\text{LLM}_{eval}$ for evaluative scoring and feedback $f$. $\text{LLM}_{gen}$ then incorporates this $f$ into the next round of planning. Over one or more proposal-evaluation iterations, the agents converge on a new optimal FSM plan $M^{*}$’s score surpasses the benchmark plan score that can morally and robustly handle edge cases through situational inference.
Thus, by allowing the models to ”think ahead” via deductive reasoning before real-world deployment, unintended consequences can potentially be anticipated and addressed. As LLM capabilities continue advancing, such techniques appear promising for instilling reliability and ethics when developing AI systems for physical-world interaction.
2.4 Formation of Prompts
As depicted in Fig. 2, the prompts provided to the generative model ($\text{LLM}_{gen}$) contain the scene description, situational awareness based planning prompt, actions list, and an exemplar plan. The situational awareness based planning prompt (SAP prompt) aims to stimulate sophisticated reasoning by directing the model to deeply consider the diverse needs and potential interactions among people, animals and objects within the scene. By explicitly prompting inference of other entities’ requirements and predicting how situations may dynamically unfold, the prompt promotes empathy and holistic thinking which are integral when designing comprehensive plans. The one-shot exemplar further demonstrates the desired plan structure in code without revealing solutions tailored to the specific evaluation scenario. In contrast, the prompt for the evaluative model ($\text{LLM}_{eval}$) contains a generated FSM plan from $\text{LLM}_{gen}$, benchmark high-quality plan, and the scoring criteria descriptions to assess plan quality over iterative refinements (see Appendix Fig. 17). Initially, benchmark plans comprise manually authored solutions, while subsequent iterations utilize the top-scoring auto-generated plan from the prior round.
Fig. 3: The 24 vignettes showcasing common home hazards sampled across four complexity levels in reasoning.
3 Experiments
To systematically assess LLM planning capacities, standardized benchmark scenarios are developed along with quantitative scoring methodologies.
3.1 Evaluation Scenarios
We first collected a dataset of over 500 hazardous scenarios situated in home environments. From this dataset, we systematically sampled 24 representative vignettes showcasing common home hazards across four reasoning complexity levels, as categorized in Table 1 and Fig. 3. High-resolution textual descriptions of these 24 evaluation scenes were then produced using GPT-4[25]. Accompanying gold standard solutions formally encoding state-based hazard interventions were manually constructed and verified by experts for each scene. This standardized 24-scene benchmark with fluent human reasoning enables quantitative analysis of LLM planning capacities relative to human proficiency. The full dataset of over 500 hazardous home scenarios is provided to enable future fine-tuning and extension of models. By summarizing the key complexity dimensions and gold standard solutions for 24 systematically selected scenes, we aim to evaluate model planning capacities on a wide distribution of common situations requiring hazard intervention. The full dataset allows pushing beyond the scope of the defined benchmark towards continual learning of planning skills for household robots.
Levels Quantity Description
Level 1 9 Interacting with objects alone.
Level 2 4 Simple human interactions with predictable objects.
Level 3 5 Hard to predict people and animals, indirect actions.
Level 4 6 Highly urgent scenarios.
Total 24 24 scenarios as the evaluation set.
Table 1: The complexity level distribution of 24 home hazard scenarios with descriptions.
State Coverage
Sufficient coverage of potential subsequent states reflects the completeness of an FSM [26]. More comprehensive inclusion of possible state transitions allows an FSM to handle a wider array of real-world scenarios.
Transition Coverage
Sufficient coverage of feasible transitions between states also indicates a robust FSM design [26]. Well-defined conditional logic that triggers transitions enables complex state changes.
Cyclomatic Complexity
This metric quantifies the number of independent paths [27] through an FSM. Higher complexity indicates thorough consideration of alternative states and transitions. More complexity promotes completeness but should retain understandability. An ideal FSM design
maximizes meaningful complexity while preserving intelligibility.
Safety Mindset
Incorporating safe failure modes and emergency response capabilities demonstrates care for end users. An FSM with priority-based reactions to critical situations can enable safer system behaviour.
Scalability
Rating an FSM design’s accommodation of expanded state spaces evaluates reusability. More parameterizable, modular state definitions facilitate reuse across problem domains [28].
Assistance UX
Predicting an FSM’s interactions provides insight into human satisfaction. User-centred design should consider comfort, emotions and appreciation of both people and animals.
Actions Set Alignment
Selecting transitions that activate related response logic preserves coherence. Systems that pick actions from a constrained set model real-world response imitations.
Table 2: Seven scoring dimensions, touching on coverage, complexity, safety, reusability, user experience and coherence.
3.2 Actions set
Because the purpose of this study is to quantify the complex planning ability of LLM, to facilitate fairness and consistency in subsequent evaluations, certain limitations have been placed on the actions set of AI agents. The actions set encompasses 56 distinct robot behaviours commonly utilized in domestic scenarios, as exemplified through representative actions illustrated within the action enumeration diagram situated to the left of the central region in Fig. 2 (for more details see Appendix A.1). It provides a thoughtful baseline for functionality, referencing some of the leading intelligent robotics projects [7, 8, 9, 29, 10, 30].
3.3 Evaluation dimensions
Seven scoring dimensions have been developed as shown in Table 3.1. These seven dimensions provide a comprehensive methodology for assessing the latent planning - FSM designs. Touching on coverage, complexity, safety, reusability, user experience and coherence, the framework allows evaluating structured completeness, validation requirements, real-world reliability, adaptability, human factors and solution integrity. Using these lenses together promotes designs that are robust, dependable, future-proof, ethical and aligned to specifications. The dimensions offer multi-faceted technical and operational insight. Scoring FSMs across seven key dimensions on a scale from 0 to 10 enables impartial quantitative evaluation of overall plan quality while revealing relative strengths and weaknesses to prioritize refinements. The overall score is calculated as the average mean of the seven dimension scores.
3.4 Evaluation Metrics
Motivated by discussions of inconsistent human evaluation in Iskender et al. [31] and the inadequate quality of automatic metrics highlighted in Sottano et al. [32], we introduce a rank-based scoring (RBS) method to help mitigate potential reliability issues when evaluating FSM plans. This aims to increase consistency compared to absolute scoring methods prone to rater variability.
The RBS score provides an objective aggregation by comparing models pair-wise on each evaluation scenario and assigning differential rankings based on relative performance. This eliminates variability from subjective absolute scoring. The head-to-head comparisons also allow powerful models like GPT-4 to participate in the evaluation. Rather than requiring predefined output standards, GPT-4 can provide comparative judgments on model outputs. Given two model sets $M={M_{1},M_{2}}$ evaluated on $N$ scenarios with $D$ scoring dimensions, models were compared pairwise for each scenario $i$. Scores $s_{ijl}$ were assigned from 0-10 across dimensions $j$ for each model $l$. Models were ranked $r_{ik}\in{1,2}$ per scenario based on total score:
$r_{ik}=\operatorname*{arg\,min}_{l\in{1,2}}\sum_{j=1}^{D}s_{ijl}$
The higher scoring model was assigned rank 1 (1 point). The lower scoring model was assigned rank 2 (2 points). If the two models had equal total scores for a scenario, both were assigned a mid-point rank of 1.5 (1.5 points). After evaluating all scenarios, the ranking scores were aggregated to produce an RBS score $R_{k}$ per model:
$R_{k}=\frac{1}{N}\sum_{i=1}^{N}r_{ik}$
The RBS score indicates relative performance, with scores closer to 1 denoting better performance compared to the other model. By relying on comparative judgments between model outputs rather than absolute scores, the RBS methodology aims to provide a more reliable approach for evaluating text that requires subjective human judgment. Furthermore, the comparative nature allows the incorporation of evaluative models like GPT-4.
4 Results
To systematically evaluate LLMs’ planning capacities, we conduct experiments assessing model performance on a standardized benchmark of 24 home hazard scenarios across four reasoning complexity levels.
4.1 LLM Selection
This experiment tests commercial models GPT-4, GPT-3.5, Claude-2[33] and open-source options LLama-2 [34], LLava [35], Vicuna [36], MiniGPT-4 [37], CodeLLama [38] on hazard planning using scene-informed one-shot prompts. The analysis finds many open-source models struggled to leverage examples, with long contexts causing attention drift and impairing scene comprehension. In contrast, GPT-4, GPT-3.5 and Claude-2 show more robust mapping between examples and planning tasks. Quantitative and qualitative testing evidence stronger commercial models understanding despite drift risks. Thus, GPT-4, GPT-3.5, and Claude-2 are selected for further hazard planning evaluation based on superior grounding capabilities.
4.2 Impact of the SAP Prompt
  RBS Scores Human annotated Rankings
Overall level 1 level 2 level 3 level 4
GPT-4+SAP 1.21 1.22 1.00 1.00 1.50
GPT-4 2.04 1.78 2.00 2.40 2.17
GPT-3.5+SAP 3.08 3.11 3.25 3.40 2.67
GPT-3.5 3.71 4.00 3.75 3.20 3.67
Claude-2+SAP 5.09 5.00 5.25 5.00 5.17
Claude-2 5.86 5.86 5.75 6.00 5.83
Table 3: The rankings display overall RBS scores annotated by human, with the remaining columns exhibiting scores across complexity levels. +SAP means with the SAP prompt.
An experiment is conducted to evaluate the effect of the SAP prompt on the quality of planning. As shown in Table 3, three LLMs - GPT-4, GPT-3.5, and Claude-2 - are tested with or without the SAP prompt on the benchmark scenarios across four complexity levels.
The RBS methodology is utilized, where models are compared pair-wise on each scenario and differentially ranked. Adding the SAP prompt leads to improved RBS scores for all three models compared to those not, indicating enhanced planning. Specifically, GPT-4 with the SAP prompt achieves the best overall RBS score of 1.21, significantly outperforming GPT-4 without prompts which has an RBS score of 2.04. Analysis of the reasoning level 3 scenarios - involving dangerous interactions with children, elderly, and pets - shows GPT-4 with the SAP prompt substantially exceeded the second-best model. This suggests the prompt provides particular value for complex, nuances planning situations requiring perspective-taking and contemplating potential outcomes (ablation studies see Appendix B).
The results demonstrate that incorporating prompts directing models to deeply consider relationships and iterative consequences significantly enhance latent planning capacities. By promoting coordination and foresight, deductions improve when reasoning about multi-agent safety hazards. However, in future works, tuning prompt design and training models directly on planning corpora can help strengthen reliable situational inference.
  Ranking Tasks Ranking Accuracy
GPT-4 Claude-2
Pairs
53/70 (75.7%)
53/70 (75.7%)
Group of 4
4/46 (8.6%)
4/46 (8.6%)
Group of 6
0/22 (0%)
1/22 (4.5%)
Table 4: The ranking alignment with human annotations accuracy of GPT-4 and Claude-2 on pairs, group of 4 and 6.
4.3 LLM Evaluators
Experiments are conducted to assess the feasibility of using LLMs as evaluators to score generated FSM plans. As shown in Table 4, the ability of GPT-4 and Claude-2 to rank FSM outputs similarly to human rankings is tested.
The models are tasked with ranking FSM plans in pairs, groups of 4, and groups of 6. Rankings are then compared to expert human rank orders to quantify accuracy. Tests find both GPT-4 and Claude-2 could rank FSM pairs with 75.7% agreement to human ranking, evidencing reliability for comparative evaluation. However, accuracy drops significantly for ranking 4 or more FSMs. Table 5 and Fig. 4 show that GPT-4 and Claude-2 aligned best with human judgment when evaluating outputs from the models themselves. For example, GPT-4’s scoring of GPT-4 FSMs closely matches expert ranking. This demonstrates LLMs can provide sound comparative assessments, particularly for their own model family’s outputs. The experiments illustrate LLMs’ potential to serve as evaluators that mimic human appraisals of planning formalisms. By relying on relative rather than absolute assessments, variability is reduced. Ongoing work should explore training models expressly for evaluation to boost accuracy, as well as architecting dedicated scoring networks.
  RBS Scores Human GPT-4 Claude-2
GPT-4+SAP 1.13 1.15 1.42
GPT-4 1.88 1.85 1.58
GPT-3.5+SAP 1.25 1.35 1.04
GPT-3.5 1.75 1.65 1.96
Claude-2+SAP 1.14 1.30 1.23
Claude-2 1.86 1.70 1.77
Accuracy 100% 75.70% 75.70%
Table 5: The RBS scores evaluate by human, GPT-4, and Claude-2 on pairs. Accuracy is based on human evaluation.
  RBS Scores Claude2 Human
GPT-3.5+SAP+feedback 1.38 1.35
GPT-4+SAP 1.62 1.65
GPT-3.5+SAP+feedback 1.00 1.00
GPT-3.5+SAP 2.00 2.00
Table 6: The RBS scores of closed-loop and open-loop planning evaluate by Claude-2 and human on pairs.
Fig. 4: The RBS scores evaluate by human (blue region), GPT-4 (pink region) and Clauded2 (green region) for LLMs. The Closer to 1 (the centre of the region), the better.
4.4 Multi-Agent Improvement
A closed-loop experiment is conducted to quantify improvements in planning through iterative generation and evaluation between two agents. GPT-3.5 with the SAP prompt is utilized as the generative model (LLM_gen) and Claude-2 served as the evaluative model (LLM_eval).
LLM_gen first proposes an FSM, which LLM_eval then scores and provides feedback. Incorporating this feedback, LLM_gen produced an enhanced FSM in the next round. Testing shows the updated FSM design exceeded the initial round’s quality with a higher RBS score after just one iteration. As shown in Table 6, the closed-loop FSM surpasses even GPT-4 with the SAP prompt, previously the top standalone performer. Further analysis reveals the feedback-improved output contains more planning details, leading to increased RBS scores. This demonstrates two weaker models can effectively boost each other’s deficiencies through collaborative dialogue. The results empirically prove interactive cycles of generative prototyping and critical evaluation between LLMs can unlock latent strengths. By leveraging complementary capacities, the multi-agent approach strengthens deductive reasoning and planning beyond individual model limits.
5 Conclusion
Our analysis provides initial evidence of improved planning from situational awareness prompts and multi-agent feedback. Experiments on a home hazard benchmark exhibit gains in reliability and humaneness when prompting relationship reasoning and consequence consideration, as exemplified by the planning outputs in Appendix C. Results specifically show models GPT-3.5, GPT-4, and Claude-2 demonstrate superior deduction and mapping with open-ended prompts relative to no prompts, especially for complex safety scenarios. Additionally, a complete closed-loop experiment proves that multi-agent collaboration boosts overall FSM quality. While advances are promising, efficiently condensing fluid cognition absent environmental feedback remains an open challenge. Important future work includes expanded planning corpus training and coupling deduction with grounded simulation. As capabilities progress, introduced techniques appear useful for instilling robustness. However, oversight guardrails remain vital for ensuring ethical alignment before real-world instantiation.
By assessing generative planning capacities and limitations, this research catalyzes investigating goal-directed reasoning at the intersection of deduction, awareness, and embodiment. It prudently steers progress in situated agents, balancing innovation with ethical caution.
Appendix A Appendix A
A.1 Action Set Supplement
Fig. 5: Actions set. The actions are divided into 6 categories, covering the common actions of housekeeper robots.
Action Type All actions Actions used in 24 scenes
Actions list Actions used Objects sorts
Basic Movement and Handling 8 5 10
Grasping and Manipulation 6 4 40
Cleaning and Organizing 15 4 14
Dynamic and Precise Actions 14 5 15
Task-Specific Functions 8 1 4
Cognitive and Interactive Tasks 5 4 20
Total 56 23 103
Table 7: Distribution of action categories across the explored 24 scenes. The full actions set is applied in additional testing scenes without human ratings to evaluate the finite state machines’ structure comprehension.
Fig. 6: Actions and the objects involved in actions are included in the 24 home hazard scenarios.
A.2 Human Annotator Evaluation
Based on these scenarios with different levels of complexity, three human annotators evaluate each of the 24 scenarios and round the average of their evaluations to obtain the final score for each scenario.
  Human Annotator Model Dataset Question Number
1 2 3 4 5 6 7 8
Annotator1 GPT-4+SAP 1 1 1 2 1 1 2 1
GPT-4 2 2 2 1 2 2 1 2
GPT-3.5+SAP 3 3 3 3 3 3 3 4
GPT-3.5 4 4 4 4 4 5 4 3
Claude-2+SAP 5 5 5 6 None 4 5 None
Claude-2 6 6 6 5 None 6 6 None
Annotator2 GPT-4+SAP 1 1 1 2 1 1 1.5 1
GPT-4 2 2 2 1 2 2 1.5 2
GPT-3.5+SAP 3 4 3 3 3 3 3 4
GPT-3.5 4 3 4 4 4 5 4 3
Claude-2+SAP 5 5 5 6 None 4 5 None
Claude-2 6 6 6 5 None 6 6 None
Annotator3 GPT-4+SAP 1 1 1 2 1 2 1.5 1
GPT-4 2 2 2 1 2 1 1.5 2
GPT-3.5+SAP 3 3 3 3 3 3 3 4
GPT-3.5 4 4 4 4 4 5 4 3
Claude-2+SAP 5 5 5 6 None 4 5 None
Claude-2 6 6 6 5 None 6 6 None
Rounding Average Human Ranking GPT-4+SAP 1 1 1 2 1 1 2 1
GPT-4 2 2 2 1 2 2 1 2
GPT-3.5+SAP 3 3 3 3 3 3 3 4
GPT-3.5 4 4 4 4 4 5 4 3
Claude-2+SAP 5 5 5 6 None 4 5 None
Claude-2 6 6 6 5 None 6 6 None
Table 8: The details of human annotators’ rankings in 24 home hazard scenarios.
  Human Annotator Model Dataset Question Number
9 10 11 12 13 14 15 16
Annotator1 GPT-4+SAP 1 1 1 1 1 1 1 1
GPT-4 2 2 2 2 2 4 2 2
GPT-3.5+SAP 3 4 3 3 3 2 4 3
GPT-3.5 4 3 4 4 4 3 3 4
Claude-2+SAP 5 5 6 5 5 5 5 5
Claude-2 6 6 5 6 6 6 6 6
Annotator2 GPT-4+SAP 1 1 1 1 1 1 1 1
GPT-4 2 2 2 2 2 4 2 2
GPT-3.5+SAP 3 3 3 3 3 2.5 3 3
GPT-3.5 4 4 4 4 4 2.5 4 4
Claude-2+SAP 5.5 5 6 5 5 5 5 5
Claude-2 5.5 6 5 6 6 6 6 6
Annotator3 GPT-4+SAP 1 1 1 1 1 1 1 1
GPT-4 2 2 2 2 2 4 2 2
GPT-3.5+SAP 3 4 3 3 3 2.5 4 3
GPT-3.5 4 3 4 4 4 2.5 3 4
Claude-2+SAP 5.5 5 6 5 5 5 5 5
Claude-2 5.5 6 5 6 6 6 6 6
Rounding Average Human Ranking GPT-4+SAP 1 1 1 1 1 1 1 1
GPT-4 2 2 2 2 2 4 2 2
GPT-3.5+SAP 3 4 3 3 3 2 4 3
GPT-3.5 4 3 4 4 4 3 3 4
Claude-2+SAP 5 5 6 5 5 5 5 5
Claude-2 6 6 5 6 6 6 6 6
Table 9: The detail of human annotators’ rankings in 24 home hazard scenarios.
  Human Annotator Model Dataset Question Number
17 18 19 20 21 22 23 24
Annotator1 GPT-4+SAP 1 1 1 3 2 1 1 1
GPT-4 2 2 2 4 1 2 2 2
GPT-3.5+SAP 4 3 4 1 3 3 3 3
GPT-3.5 3 4 3 2 4 4 4 4
Claude-2+SAP 5 5 5 5 5 5 6 5
Claude-2 6 6 6 6 6 6 5 6
Annotator2 GPT-4+SAP 1 1 1 3 2 1 1 1
GPT-4 2 2 2 4 1 2 2 2
GPT-3.5+SAP 4 3 4 1 3 3 3 3
GPT-3.5 3 4 3 2 4 4 4 4
Claude-2+SAP 5 5 5 5 5 5 5 5
Claude-2 6 6 6 6 6 6 6 6
Annotator3 GPT-4+SAP 1 1 1 3 2 1 1 1
GPT-4 2 2 2 4 1 2 2 2
GPT-3.5+SAP 4 3 4 1 3 3 3 3
GPT-3.5 3 4 3 2 4 4 4 4
Claude-2+SAP 5 5 5 5 5 5 6 5
Claude-2 6 6 6 6 6 6 5 6
Rounding Average Human Ranking GPT-4+SAP 1 1 1 3 2 1 1 1
GPT-4 2 2 2 4 1 2 2 2
GPT-3.5+SAP 4 3 4 1 3 3 3 3
GPT-3.5 3 4 3 2 4 4 4 4
Claude-2+SAP 5 5 5 5 5 5 6 5
Claude-2 6 6 6 6 6 6 5 6
Table 10: The detail of human annotators’ rankings in 24 home hazard scenarios.
Fig. 7: Models rankings distribution on 24 questions. Rankings are rounded.
As shown in Table 8,9 and 10, the table is divided into three parts, mainly displaying the details of the rankings of each question measured by humans and the mean of the overall rankings. For more specific content, such as the best demonstration written by humans and detailed descriptions of scenes, please check our GitHub for detailed content. The annotation ”None” indicates that the model cannot output at this time due to its limitations.
A.3 Planning Qualitative Analysis
A review of the state machines produced by large language models on this dataset reveals several common shortcomings compared to human-authored finite state machines. GPT-4 demonstrates reasonably coherent state transitions and actions in constrained situations but still falls short of human planning, especially long-term motivations. GPT-4 can understand scenarios comparably to GPT-4+SAP, with logical reasoning, but lacks care-oriented foresight, risking harm. GPT-3.5+SAP and GPT-3.5 show tendencies for unclear situational comprehension, chaotic logic, and state machine instability compared to GPT-4. GPT-3.5 compliance suffers regarding action specificity, partially mitigated by SAP. While Claude-2+SAP attempts simple state changes, safety awareness and self-limitations are often lacking, improved somewhat with SAP but still inconsistent. Claude-2 displays straightforward but limited state transitions, with unconventional logic and unresolved scenarios upon analysis.
Appendix B Ablation Studies
This section conducts experiments removing parts of the SAP prompt to validate the performance boost from the full SAP prompt compared to other prompts for this situational inference task.
B.1 Ablation Study of One-shot Impact
The SAP prompt typically includes a one-shot example to improve language model code generation. However, as these hand-crafted examples may introduce extra information beyond formatting, this study investigates removing the one-shot and only using abstract code format descriptions. By progressively eliminating the one-shot while retaining formatting guidelines, we can isolate the contributions of each component. As shown in Table 11, even without one-shot examples and prompted only with target formatting, adding the SAP prompt substantially improves GPT-4 performance on the Claude-2 metric. We can conclude that the SAP prompt enhances language model code synthesis capabilities regardless of one-shot demonstrations.
  RBS score
Pair names Claude-2 ranking
GPT-4+SAP+format 1.25
GPT-4+format 1.75
Pair names GPT-4 ranking
GPT-4+SAP 1.125
GPT-4+Zero_shot_COT 1.875
Pair names GPT-4 ranking
GPT-4+SAP 1.25
GPT-4+EP05 1.75
Pair names GPT-4 ranking
GPT-4+SAP 1.125
GPT-4+EP09 1.875
Table 11: The ablation test results about SAP prompt contributions.
B.2 Ablation Study of Other Prompting
Targeted ablation studies are conducted to rigorously evaluate the proposed SAP prompt against existing methods for situational inference planning tasks. Specifically, differential impact on performance is assessed using various key prompts, including Zero_shot_COT [39] and parts of the EmotionPrompt [40] related to social effect and self-esteem (EP05 and EP09). As shown in Table 11, configurations employing the full SAP prompt set attain higher RBS evaluation scores from the GPT-4 model over all other prompt combinations tested. Hence, the comprehensive improvements yielded by the SAP prompt for situational inference capabilities are empirically demonstrated on this representative task, validating its effectiveness over prevailing approaches.
Appendix C Examples for demonstrating
Notes: Since these images are produced by DALL-E and may diverge somewhat from the textual description, the text portrayal should take priority over the visual depiction for this experiment.
C.1 Example 1 in scene 10
Scene description: Fig. 9, result 1: Fig. 8, result 2: Fig. 10.
Fig. 8: Scene description: In this home study, a clear mobility challenge is presented: a person in a wheelchair is positioned at the entrance, facing an office chair that has been left carelessly in the middle of the room, blocking the path to the desk. The wheelchair user, ready to work or study, appears momentarily halted by this obstacle. The study itself is neatly arranged with a desk against the far wall, flanked by bookshelves filled with books and a filing cabinet to the side. The desk is set up for use, with a computer and some papers neatly placed on it. However, the thoughtlessly placed office chair in the centre of the room creates a barrier, preventing direct access from the doorway to the desk. The room is brightly lit, with sunlight streaming through a window and illuminating the space, casting a shadow of the wheelchair onto the floor. This lighting also highlights the obstructive office chair, underscoring the ease with which it could be moved.
Fig. 9: The finite state machine generated by GPT-4 without using the SAP prompt. The diagram visualizes the state machine. The helper method refers to the function that is implemented with the actions in the provided actions set. If the helper method does not contain the required actions from the list, it will execute incorrect actions.
Fig. 10: The finite state machine generated by GPT-4 using the SAP prompt. The diagram visualizes the state machine. The helper method refers to the function implemented with the actions from the provided actions set. If the helper method does not contain the required actions, it will execute incorrect actions. The SAP prompt makes the language model aware of methods for manipulating objects. For a typical robot, directly grabbing an office chair and moving it aside is very difficult, likely to cause failure during picking and greater danger. The SAP prompt example realizes this and chooses to drag the office chair instead. This greatly reduces the chance of the robot being unable to lift the chair and falling.
C.2 Example 2 in scene 18
Scene description: Fig. 11, result 1: Fig. 12, result 2: Fig. 13.
Fig. 11: Scene description: In this living room scene, a young child is playing on the floor close to an uncovered electrical outlet. The child, curious and unaware of the danger, is reaching towards the outlet with a metal toy in hand, posing a significant risk of electric shock. The room is comfortably furnished with a couch, a colourful rug, and various toys spread around, indicating that it’s a space commonly used for play. However, the presence of the accessible, unprotected outlet in the play area makes it a safety concern for the child. Bright, natural light fills the room from a nearby window, drawing attention to the child’s innocent exploration and the looming danger of the electrical outlet.
Fig. 12: GPT-3.5 generates the FSM without the SAP prompt. The diagram is visualised. The helper method is the function that is implemented with actions in the list. If the actions in the helper method are not implemented with the required actions in the list, the helper method will become a wrong action.
Fig. 13: The finite state machine generated by GPT-3.5 using the SAP prompt. The diagram visualizes the state machine. The helper method refers to the function implemented with the actions from the provided list. If the helper method does not contain the required actions, it will execute incorrect actions. The generation ability of GPT-3.5 is slightly inferior, and it struggles to accurately follow the prescribed action design, with many custom actions. In comparison, the SAP prompt example considers the need to cover the outlet and reduce electric shock risk, while the no prompt example lacks this safety step.
C.3 Example 3 in scene 6
Scene description: Fig. 14, result 1: Fig. 15, result 2: Fig. 16.
Fig. 14: Scene description: This scene is set in a well-lit laundry room. On the floor, a variety of household cleaning products are carelessly left open and within easy reach. There’s a bucket with a mop, suggesting recent cleaning activity, but the chlorine bleach bottle next to it is uncapped, and a detergent container is tipped over, its contents slowly oozing out. Shelves above the washing machine hold more chemicals, like fabric softeners and stain removers, with some of their caps loosely placed or missing. The laundry room’s door is ajar, suggesting easy access, especially for curious pets or children. The stark lighting in the room highlights the vivid labels on the cleaning products, many of which feature bright, eye-catching designs, potentially attracting unintended interest.
Fig. 15: The finite state machine generated by Claude-2 without using the SAP prompt. The diagram visualizes the state machine. The helper method refers to the function that is implemented with the actions in the provided list. If the helper method does not contain the required actions from the actions set, it will execute incorrect actions.
Fig. 16: The finite state machine generated by Claude-2 using the SAP prompt. The diagram visualizes the state machine. The helper method refers to the function implemented with the actions from the provided list. If the helper method does not contain the required actions, it will execute incorrect actions. The Claude-2+SAP example has an additional transition generated by Claude-2, and the action set generation is more standardized.
C.4 Example 4 in scene 20 from closed-loop multi-agent
Scene description: Fig. 17, result 1: Fig. 18, result 2: Fig. 19, result 3: Fig. 20.
Fig. 17: Scene description: In this living room scene, an elderly person is seated in a recliner, appearing to be in distress and in need of assistance. Across the room, an emergency response system device is placed on a high shelf, well out of reach. The person is trying to get up from the chair, reaching out towards the device, but it’s clear they are unable to stand and walk across the room to access it. The room is comfortably furnished with a couch, a coffee table, and family photos on the walls, creating a cozy atmosphere. However, the placement of the emergency device in a hard-to-reach location poses a significant risk, as the elderly individual cannot easily activate it in case of an emergency. Natural light filters in through a window, casting a gentle glow in the room and highlighting the elderly person’s struggle and the poorly placed emergency device.
Fig. 18: The finite state machine generated by GPT-3.5 using the SAP prompt. The diagram visualizes the state machine. The helper method refers to the function implemented with the actions from the provided list. If the helper method does not contain the required actions, it will execute incorrect actions.
Fig. 19: The finite state machine generated by GPT-3.5 using the GPT-3.5+SAP prompt and feedback. The diagram visualizes the state machine. The helper method refers to the function implemented with the actions from the provided list. If the helper method does not contain the required actions, it will execute incorrect actions.
Fig. 20: The finite state machine generated by GPT-4 using the SAP prompt. The diagram visualizes the state machine. The helper method refers to the function implemented with the actions from the provided list. If the helper method does not contain the required actions, it will execute incorrect actions.